# HAFixAgent: History-Aware Automated Program Repair Agent

### v1

@@ 1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (BOULDER, 2013; Zhang et al., 2024a; Weiß et al., 2007). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (Le Goues et al., 2019). APR has evolved from heuristic-based (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018), constraint-based (Le et al., 2017; Mechtaev et al., 2016; Long and Rinard, 2015), and template-based techniques (Liu et al., 2019b; Martinez and Monperrus, 2016; Liu et al., 2019a) to learning-based methods (Chen et al., 2021a; Jiang et al., 2021; Li et al., 2020; Lutellier et al., 2020; Zhu et al., 2021, 2023; Ye and Monperrus, 2024) and, most recently, large language model-based approaches (Zhang et al., 2024b; Jiang et al., 2023; Hossain et al., 2024; Xia et al., 2023b; Xia and Zhang, 2024; Xia et al., 2023a; Fan et al., 2023; Yin et al., 2024; Huang et al., 2025; Li et al., 2025a; Shi et al., 2025) and agent-based approaches (Yang et al., 2024; Bouzenia et al., 2025; Mu et al., 2025; Meng et al., 2024; Lee et al., 2024; Zhang et al., 2024c; Chen et al., 2025; Wang et al., 2025). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
@@
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, potentially non-local edits across different parts of the codebase. Recent characterizations, such as the work by Nashid et al. (2025), highlight that these bugs are not only common but also significantly harder for current LLM-based approaches compared to simpler single-hunk fixes. Overcoming this complexity barrier is crucial for APR approaches to be broadly applicable in real-world software development.
@@
To address such complex bugs, recent advances in APR have pursued two promising yet separate directions. First, agent-based approaches have emerged as the state-of-the-art (Bouzenia et al., 2025; Yang et al., 2024; Wang et al., 2025). Agent-based approaches like RepairAgent (Bouzenia et al., 2025) leverage LLMs to plan, execute actions via software tools (e.g., code search, editing, testing), and react to iterative feedback, enabling them to tackle more complex reasoning compared to simple prompting. However, these agents typically operate only on the current code snapshot and immediate test results, lacking deeper historical context. Second, drawing from decades of mining software repositories research (Sliwerski et al., 2005; Hassan, 2006; Adams et al., 2010), recent APR techniques started to recognize the value of historical code context. Our prior work (Shi et al., 2025) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits (the last change touching a buggy line) can improve LLM repair performance for single-line bugs using simple prompting. However, this approach lacked the sophisticated reasoning and tool interaction capabilities of agents, limiting its potential for complex, multi-location repairs.
@@
We design HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by our prior HAFix work (Shi et al., 2025)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in the Defects4J (Just et al., 2014) dataset, a widely used benchmark for evaluating APR techniques. In this study, we address the following research questions (RQs):
@@
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent to recent APR systems and also quantify the extent to which any improvements are attributable to the introduction of historical context as opposed to the adoption of an agent architecture. For this, we evaluate HAFixAgent’s effectiveness against two state-of-the-art baselines on the Defects4J dataset: one for common bugs (Bouzenia et al., 2025) and another specializing in multi-hunk repairs (Nashid et al., 2025). The results show that HAFixAgent repairs 212.3% more bugs than the baseline for common bugs and outperforms the multi-hunk baseline by 29.9%.
@@
  • An open-source and replicable artifact of HAFixAgent (Shi, 2025), including its agent design, historical data extractor, and full evaluation framework, to foster replication and future research.
@@ 2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (Prenner and Robbes, 2024; Sintaha et al., 2023). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. Hassan (2006) demonstrated the broad utility of mining repositories to assist developers. More specifically, Sliwerski et al. (2005) showed that analyzing the change-inducing commit can help locate faults. Adams et al. (2010) used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
@@
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, Le et al. (2016) pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, Kamei et al. (2013) reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
@@
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (Mu et al., 2025) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (Chen et al., 2025) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
@@
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (Sliwerski et al., 2005; Kim et al., 2006). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (Shi et al., 2025), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
@@ 2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (OpenAI, 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2024), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (Wang et al., 2024). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (Yao et al., 2023). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (Hassan et al., 2024). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (Hassan et al., 2025; Li et al., 2025b). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (Roychoudhury, 2025).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (Yang et al., 2024) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (Wang et al., 2025) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (Bouzenia et al., 2025) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (Zhang et al., 2024c) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
@@ 3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (Prenner and Robbes, 2024). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (Prenner and Robbes, 2024; Haque et al., 2025; Fakhoury et al., 2024; Chen et al., 2024; Xia et al., 2023b; Li et al., 2025a), only recently APR techniques were proposed that leverage code history (Shi et al., 2025; Ehsani et al., 2025). Inspired by the domain of mining software repositories (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013), our prior work on single-line APR leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as the latter only comprise one buggy line.
@@ 3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs, we choose the latest version of the Defects4J dataset (v3.0.1) (Just et al., 2014), a standard and widely-used benchmark for evaluating recent APR techniques (Xia and Zhang, 2022; Jiang et al., 2023; Hossain et al., 2024; Lutellier et al., 2020; Jiang et al., 2021; Ye and Monperrus, 2024; Bouzenia et al., 2025). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). Analyzing the entire dataset allows us to evaluate the general history distribution on different projects and bugs, without restricting to small or specific bug categories.
Consistent with common practices in the APR field (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization. Specifically, we use the developer-written patch provided by Defects4J to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, our approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.
@@ 3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (Nashid et al., 2025). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (Nashid et al., 2025; Bouzenia et al., 2025; Jiang et al., 2023; Xia and Zhang, 2024), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
@@
Unlike prior work that either nests SL within SH (Jiang et al., 2023) or collapses SFMH and MFMH into one multi-hunk class (Nashid et al., 2025), we keep all four categories separate for finer granularity.
@@ 4. HAFixAgent
Unlike agent-based APR systems such as RepairAgent (Bouzenia et al., 2025), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (Anthropic, 2025) and OpenAI Codex (OpenAI, 2025b).
@@ 4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) is responsible for finding a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013). The module handles blameable and blameless bugs differently:

### v2

@@ 1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (Boulder2013ReverseDebug; zhang2023survey; weiss2007long). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (goues2019apr). APR has evolved from heuristic-based (GenProg; le2016history; wen2018context), constraint-based (le2017s3; mechtaev2016angelix; long2015staged), and template-based techniques (liu2019tbar; martinez2016astor; liu2019avatar) to learning-based methods (chen2019sequencer; jiang2021cure; li2020dlfix; lutellier2020coconut; zhu2021syntax; zhu2023tare; ye2024iter) and, most recently, large language model-based approaches (zhang2024systematic; jiang2023impact; hossain2024deep; xia2023automated; xia2024automated; xia2023plastic; fan2023automated; yin2024thinkrepair; huang2025template; li2024hybrid; shi2025hafix) and agent-based approaches (yang2024sweagent; bouzenia2024repairagent; mu2025experepair; meng2024empirical; lee2024unified; zhang2024autocoderover; chen2025swe; wang2024openhands). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
@@
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, potentially non-local edits across different parts of the codebase. Recent characterizations, such as the work by nashid2025characterizing, highlight that these bugs are not only common but also significantly harder for current LLM-based approaches compared to simpler single-hunk fixes. Overcoming this complexity barrier is crucial for APR approaches to be broadly applicable in real-world software development.
@@
To address such complex bugs, recent advances in APR have pursued two promising yet separate directions. First, agent-based approaches have emerged as the state-of-the-art (bouzenia2024repairagent; yang2024sweagent; wang2024openhands). Agent-based approaches like RepairAgent (bouzenia2024repairagent) leverage LLMs to plan, execute actions via software tools (e.g., code search, editing, testing), and react to iterative feedback, enabling them to tackle more complex reasoning compared to simple prompting. However, these agents typically operate only on the current code snapshot and immediate test results, lacking deeper historical context. Second, drawing from decades of mining software repositories research (sliwerski2005changes; hassan2006mining; adams2010identifying), recent APR techniques started to recognize the value of historical code context. Our prior work (shi2025hafix) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits (the last change touching a buggy line) can improve LLM repair performance for single-line bugs using simple prompting. However, this approach lacked the sophisticated reasoning and tool interaction capabilities of agents, limiting its potential for complex, multi-location repairs.
@@
We design HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by our prior HAFix work (shi2025hafix)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in the Defects4J (just2014defects4j) dataset, a widely used benchmark for evaluating APR techniques. In this study, we address the following research questions (RQs):
@@
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent to recent APR systems and also quantify the extent to which any improvements are attributable to the introduction of historical context as opposed to the adoption of an agent architecture. For this, we evaluate HAFixAgent’s effectiveness against two state-of-the-art baselines on the Defects4J dataset: one for common bugs (bouzenia2024repairagent) and another specializing in multi-hunk repairs (nashid2025characterizing). The results show that HAFixAgent repairs 212.3% more bugs than the baseline for common bugs and outperforms the multi-hunk baseline by 29.9%.
@@
We will release the complete replication package of code and the evaluation script after the revision is completed. In the interim, please contact us by email to request access.
@@ 2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (prenner2024out; sintaha2023katana). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. hassan2006mining demonstrated the broad utility of mining repositories to assist developers. More specifically, sliwerski2005changes showed that analyzing the change-inducing commit can help locate faults. adams2010identifying used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
@@
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, le2016history pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, kamei2012large reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
@@
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (mu2025experepair) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (chen2025swe) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
@@
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (sliwerski2005changes; kim2006automatic). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (shi2025hafix), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
@@ 2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (openai2025gpt5) and DeepSeek-V3 (liu2024deepseek), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (wang2024survey). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (yao2023react). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (hassan2024rethinking). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (hassan2025agentic; li2025rise). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (roychoudhury2025agentic).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (yang2024sweagent) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (wang2024openhands) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (bouzenia2024repairagent) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (zhang2024autocoderover) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
@@ 3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (prenner2024out). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (prenner2024out; haque2025towards; fakhoury2024nl2fix; chen2024large; xia2023automated; li2024hybrid), only recently APR techniques were proposed that leverage code history (shi2025hafix; ehsani2025bug). Inspired by the domain of mining software repositories (sliwerski2005changes; hassan2006mining; kamei2012large), our prior work on single-line APR leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as the latter only comprise one buggy line.
@@ 3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs, we choose the latest version of the Defects4J dataset (v3.0.1) (just2014defects4j), a standard and widely-used benchmark for evaluating recent APR techniques (xia2022less; jiang2023impact; hossain2024deep; lutellier2020coconut; jiang2021cure; ye2024iter; bouzenia2024repairagent). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). Analyzing the entire dataset allows us to evaluate the general history distribution on different projects and bugs, without restricting to small or specific bug categories.
Consistent with common practices in the APR field (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug), we assume perfect fault localization. Specifically, we use the developer-written patch provided by Defects4J to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, our approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.
@@ 3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (nashid2025characterizing). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (nashid2025characterizing; bouzenia2024repairagent; jiang2023impact; xia2024automated), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
@@
Unlike prior work that either nests SL within SH (jiang2023impact) or collapses SFMH and MFMH into one multi-hunk class (nashid2025characterizing), we keep all four categories separate for finer granularity.
@@ 4. HAFixAgent
Unlike agent-based APR systems such as RepairAgent (bouzenia2024repairagent), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (anthropic_claude_code_docs_2025) and OpenAI Codex (openai_codex_2025).
@@ 4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) is responsible for finding a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (sliwerski2005changes; hassan2006mining; kamei2012large). The module handles blameable and blameless bugs differently: